Hi! I have to write an algebraic equation with clearly defined variables to represent the cost of each option. I also have to be able to graph the equation in excel. Here is my word problem:
A working parent is looking for child care for their child. Day care A (a home-based option) charges a flat rate of $5 per hour. Day care B (a center based option) charges a fixed rate of $100 per week and $10.00 per hour for any time over 40 hours per week. If the parent works 50 hours per week which option is the best value for the money?
Can you help me write the appropriate equation for this problem. Thanks so much .

Answer by stanbon(75887)

A working parent is looking for child care for their child.
Day care A (a home-based option) charges a flat rate of $5 per hour.
--------------------
A(h = 5h where h is number of hours
----------------------
Day care B (a center based option) charges a fixed rate of $100 per week and $10.00 per hour for any time over 40 hours per week.
------
If the parent works 50 hours per week which option is the best value for the money?
Cost for 50 hrs in one week:
A(50) = 5*50 = $250.00
B(50) = 100 + 10(50-40) = 100 + 100 = $200.00
---
Cheers,
Stan H.